OpenStudy (anonymous):

Edith buys last years best selling novel in hardcover, for $19.55. This is with a 15% discount from the original price. What was the original price of the novel?

OpenStudy (anonymous):

23.00 as 19.55/85 =0.23 0.23*100= 23.00

OpenStudy (anonymous):

19.55=85%*x 19.55=0.85x x=23 origin price =23 x is the original price

OpenStudy (anonymous):

sorry i will elaborate if 19.55 is after a 15% discount then it is 85% of the original value

OpenStudy (anonymous):

$19.55 + 15% = $22.48

OpenStudy (anonymous):

so $23.00 is the correct answer?

OpenStudy (anonymous):

$19.55 = 85% of original price 19.55 /85 = 1% = 0.23 now multiply this by 100 to get to the original price 0.23 x 100 = 23